OpenStudy (anonymous):

What is the slope of the line whose equation is 5y + 6x - 2 = 0? slope = 6 slope = -6/5 slope = 5/6

OpenStudy (anonymous):

get everything in terms of y 5y + 6x - 2 = 0 5y + 6x = 2 5y = -6x + 2 y = -6/5x + 2/5 this is in the form y = mx + b where m is the slope and b is the y-intercept therefore m = -6/5

OpenStudy (anonymous):

thanks for the explain to

OpenStudy (skullpatrol):

Of course you are going to thank him for doing all the work for you.

OpenStudy (anonymous):

i dont need any input from you ok

OpenStudy (anonymous):

he explained it and i can learn from that could you not see that or did you just decide oh this is a great day to be an as* about eveyrthing

OpenStudy (anonymous):

Why does everyone assume I'm a guy??